OpenStudy (anonymous):

6(2p+q)^2-5(2p+q)-25

OpenStudy (aum):

6(2p+q)^2 - 5(2p+q) - 25 Let t = (2p+q). Then, the above expression becomes: 6t^2 - 5t - 25 Use the AC method to factor: a = 6, b = -5, c = -25 a * c = 6 * (-25) = -150. What two numbers multiply to -150 and add to -5? -15 and 10 6t^2 - 5t - 25 = 6t^2 - 15t + 10t - 25 = 3t(2t - 5) + 5(2t - 5) = (2t - 5)(3t + 5) Replace t by (2p+q) ( 2(2p+q) - 5 ) * ( 3(2p+q) + 5 ) = (4p + 2q - 5)(6p + 3q + 5)